OpenStudy (anonymous):

The ratio between the radii of two spheres is 9:2 What is the ratio of their volumes?

OpenStudy (anonymous):

cube the ratios

OpenStudy (anonymous):

is it 729:8 ?

OpenStudy (anonymous):

r:R is ratio between radii then r^3:R^3 is ratio between volume

OpenStudy (anonymous):

so am i correct then?

OpenStudy (anonymous):

\[V1/V2 = (4*\pi*r1^3)/(4*\pi*r2^3)\]

OpenStudy (anonymous):

Volume is proportional to the cube of the radius

OpenStudy (anonymous):

im confused am i wrong then?

OpenStudy (kropot72):

@ineedhelponmyhw Your answer 729/8 is correct.

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

Answer is correct.....i was just giving u the concept @ineedhelponmyhw

OpenStudy (kropot72):

You're welcome :)